OpenStudy (anonymous):

What is the approximate area of a sector with radius equal to 8 and measure of arc equal to 300°? 3.3π 13.3π 53.3π

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

13.3?

OpenStudy (anonymous):

@geerky42

TheSmartOne (thesmartone):

Nope.

OpenStudy (anonymous):

all right thnx

TheSmartOne (thesmartone):

\(\sf\Large \pi r^2 \times \frac{central ~angle}{360}\)